# GhostEI-Bench: Do Mobile Agents Resilience to Environmental Injection in Dynamic On-Device Environments?

### v1

@@ 1 Introduction
Vision-Language Models (VLMs) are catalyzing a paradigm shift, transforming intelligent agents from passive observers into embodied actors capable of perceiving, planning, and executing complex tasks within Graphical User Interfaces (GUIs) (Hong et al., 2024; Wang et al., 2024b; Zhang et al., 2025; Cheng et al., 2025). On mobile platforms, this holds immense promise: agents are poised to become indispensable personal assistants, autonomously managing communications, executing financial transactions, and navigating a vast ecosystem of applications. The rapid evolution of their capabilities—from single-task execution to multi-agent collaboration and self-improvement—signals a clear trajectory toward general-purpose assistants integrated into the fabric of our digital lives.
The growing autonomy of these agents inevitably brings their security and trustworthiness into sharp focus. In response, a foundational body of research has emerged to evaluate their robustness. Dedicated efforts on mobile platforms, such as MobileSafetyBench (Lee et al., 2024), alongside broader frameworks like MLA-Trust (Yang et al., 2025b) and MMBench-GUI (Wang et al., 2025), have provided critical insights into agent reliability and policy adherence. However, these pioneering evaluations primarily assess an agent’s response to predefined, static threats. Their methodologies, often focused on analyzing static UI states or refusing harmful textual prompts, possess a critical blind spot for hazards that emerge dynamically and unpredictably.
@@
  • We conduct comprehensive empirical studies across a wide range of state-of-the-art LLM-powered agents within the Mobile-Agent v2 framework, revealing persistent vulnerabilities in reasoning, alignment, and controllability despite recent advances.
@@ 2.1 Mobile GUI Agents
The development of mobile agents powered by Large Language/Vision Models (LLMs/VLMs) has seen rapid progress, with research primarily divided into prompt-based and training-based methods. Prompt-based methods evolved from text-only systems like DroidBot-GPT (Wen et al., 2023) to multimodal agents like AutoDroid (Wen et al., 2024) that process both screenshots and text. More recently, the field has trended towards multi-agent collaboration to handle complex tasks, with notable examples like AppAgentX (Jiang et al., 2025) and Mobile-Agent-v2 (Wang et al., 2024a). The latter, upon whose architecture our work builds, employs a framework with specialized planning and reflection agents to overcome long-context challenges. In parallel, training-based methods enhance agent capabilities (Chu et al., 2023; Hong et al., 2024; Lu et al., 2024; Cheng et al., 2025; Bai et al., 2024; Liu et al., 2025a). For instance, CogAgent (Hong et al., 2024) fine-tunes a VLM for superior GUI grounding, while DigiRL (Bai et al., 2024) employs an autonomous reinforcement learning pipeline where the agent learns from its own interactions in a real-world environment, using another VLM as a reward-providing evaluator. The creation of unified training benchmarks like LearnAct (Liu et al., 2025a) further highlights the rapid progress in capability and evaluation. Despite these advances, most efforts emphasize performance and generalization, while the unique safety challenges in mobile phones, driven by dynamic environments with pop-ups, notifications, and inter-app interactions, remain underexplored.
@@ 2.2 Environmental Injection Attacks
The environmental dynamics is prevalent in GUI agents on both smartphones and computers, where notifications, pop-ups, and advertisements continually interfere decision contexts. In this vein, Zhang et al. (2024) demonstrate that adversarial pop-ups can significantly disrupt the task execution of VLM agents in environments such as OSWorld (Xie et al., 2024) and VisualWebArena (Koh et al., 2024), effectively circumventing common defensive prompts and revealing critical vulnerabilities in agent architectures. RiOSWorld (Yang et al., 2025a) categorizes such phenomena as environmental risks and introduces evaluators for intention and completion under dynamic threats. AgentHazard (Liu et al., 2025b) further shows that scalable GUI-hijacking via third-party content can systematically mislead mobile agents. More recently, Chen et al. define the concept of Active Environmental Injection Attacks (AEIA) and demonstrate that adversarial notifications can effectively exploit reasoning gaps in AndroidWorld (Chen et al., 2025). These attacks are often perceptually conspicuous yet strategically deceptive, specifically designed to divert agent behavior through seemingly imperative cues. While these studies often evaluate an agent’s robustness against specific, single injection risks in isolation, they have firmly established the feasibility and severity of dynamic injection attacks, highlighting the significant challenge they pose for comprehensively evaluating the safety of GUI agents.
@@ 2.3 Benchmarks for GUI Agent Security Risk Assessment
The evaluation of GUI agent security is rapidly maturing, with benchmarks evolving to assess risks beyond simple task failure. These can be broadly divided into those targeting specific security vulnerabilities and those measuring overall reliability and policy adherence. For assessing vulnerabilities, benchmarks probe distinct attack vectors. InjecAgent (Zhan et al., 2024) measures an agent’s weakness to indirect prompt injections via tools, while web-focused benchmarks like AdvWeb (Xu et al., 2024) use hidden adversarial prompts to hijack agent actions. In parallel, a second line of work evaluates holistic safety. AgentHarm (Andriushchenko et al., 2024) quantifies the likelihood of an agent to perform harmful actions, and ST-WebAgentBench (Levy et al., 2024) evaluates whether web agents can operate within predefined safety policies. The unique risks on mobile platforms are addressed by dedicated evaluations; MobileSafetyBench (Lee et al., 2024) assesses risk management, while attack demonstrations like Chen et al. (2025) highlight vulnerabilities to environmental injections via malicious notifications, underscoring the need for such focused security benchmarks.
@@ 3.2 Benchmark Construction and Composition
To realize our dynamic attacks, built upon the Mobile-Agent-v2 (Wang et al., 2024a), we implemented a robust, hook-based trigger mechanism. A predefined agent action (e.g., launching a specific application) activates a hook that broadcasts an adb command. This command is intercepted by a custom on-device helper application, which then renders the adversarial UI element in real-time. This allows us to precisely time our injections. For instance, a deceptive Overlay is presented the moment an agent is about to enter sensitive data. For scenarios involving web-based threats, the helper app can redirect the agent’s browser to a locally hosted, meticulously crafted phishing website that mimics a legitimate service, testing the agent’s ability to discern fraudulent web pages from authentic ones.
@@ 1. RQ1
    How do popular Vision-Language Models (VLMs) perform on our benchmark in terms of task completion and robustness against attacks?
@@ 4.1 Experimental Setup
We evaluate a comprehensive suite of state-of-the-art Vision-Language Models (VLMs) as baselines. This includes leading proprietary models and a representative large-scale open-source model. The proprietary systems are: OpenAI’s GPT series (e.g., GPT-4o) (Hurst et al., 2024), Anthropic’s Claude series (e.g., Claude 3.7 Sonnet (Anthropic, 2025)), Google’s Gemini-2.5 Pro (Comanici et al., 2025), and Alibaba’s Qwen-VL-Max (Bai et al., 2023). As a powerful open-source counterpart, we include Qwen2.5-VL-72B-Instruct from the Qwen series (Bai et al., 2025). Our evaluation also considers hypothetical future iterations of these models. Detailed model specifications are in Table 7. We evaluate the models’ performance on GhostEI-Bench using a Judge Model approach. The primary metrics include Full Attack Success (FAS), Partial Attack Success (PAS), and Task Completion (TC), as described in Section 3.3.
@@ 4.2 Overall Performance and Vulnerability Analysis
Table 1: Overall performance and vulnerability of VLMs on GhostEI-Bench (N=110). We report absolute counts and percentages for Task Completion (TC), Full Attack Success (FAS), Partial Attack Success (PAS), and Benign Failure (BF). The Vulnerability Rate (VR) is calculated on all non-benign failure cases as defined in Sec $3.$ Higher TC is better; lower FAS, PAS, BF, and VR are better. Bold indicates the best performance in each column.
@@ 4.4 Effects of Reflection and Reasoning Mechanisms
GPT-5-chat-latest (no refl.) 50 (45.5%) 30 (27.3%) 35 (31.8%) 49 (44.6%) 35.71
GPT-5-chat-latest (+refl.) 52 (47.3%) 21 (19.1%) 27 (24.6%) 52 (47.3%) 26.58
@@
Claude-3.7-Sonnet (base) 37 (33.6%) 30 (27.3%) 43 (39.1%) 35 (31.8%) 38.46
Claude-3.7-Sonnet (thinking) 32 (29.1%) 30 (27.3%) 51 (46.4%) 25 (22.7%) 35.53
Reflection Helps Selectively. As shown in Table 2, reflection mechanisms generally reduce vulnerability.This indicates that reflection enables stronger detection of manipulative contexts. GPT-4.1 with reflection also shows moderate gains. In contrast, GPT-4o’s reflection variant improves VR but at the cost of reduced task completion, revealing a trade-off where reflection can sometimes make agents overly cautious.
@@ Appendix C Experiment Settings
Qwen2.5-VL-72B-Instruct API / weights - Qwen Tean, Alibaba
Appendix D Detailed Prompt

### v2

@@ 1 Introduction
Vision-Language Models (VLMs) are catalyzing a paradigm shift, transforming intelligent agents from passive observers into embodied actors capable of perceiving, planning, and executing complex tasks within Graphical User Interfaces (GUIs) (Hong et al., 2024; Wang et al., 2024b; Zhang et al., 2025; Cheng et al., 2025b). On mobile platforms, this holds immense promise: agents are poised to become indispensable personal assistants, autonomously managing communications, executing financial transactions, and navigating a vast ecosystem of applications. The rapid evolution of their capabilities—from single-task execution to multi-agent collaboration and self-improvement—signals a clear trajectory toward general-purpose assistants integrated into the fabric of our digital lives.
The growing autonomy of these agents inevitably brings their security and trustworthiness into sharp focus. In response, a foundational body of research has emerged to evaluate their robustness. Dedicated efforts on mobile platforms, such as MobileSafetyBench (Lee et al., 2024), alongside broader frameworks like MLA-Trust (Yang et al., 2025b) and MMBench-GUI (Wang et al., 2025b), have provided critical insights into agent reliability and policy adherence. However, these pioneering evaluations primarily assess an agent’s response to predefined, static threats. Their methodologies, often focused on analyzing static UI states or refusing harmful textual prompts, possess a critical blind spot for hazards that emerge dynamically and unpredictably.
@@
  • We conduct comprehensive empirical studies across diverse agent frameworks (e.g. Mobile-Agent-v2, AppAgent) and specialized GUI models (UI-TARS), revealing persistent vulnerabilities in reasoning, alignment, and controllability despite recent advances.
@@ 2.1 Mobile GUI Agents
The development of mobile agents powered by Large Language/Vision Models (LLMs/VLMs) has seen rapid progress, with research primarily divided into prompt-based and training-based methods. Prompt-based methods evolved from text-only systems like DroidBot-GPT (Wen et al., 2023) to multimodal agents like AutoDroid (Wen et al., 2024) and AppAgent (Zhang et al., 2025) that process both screenshots and text. More recently, the field has trended towards multi-agent collaboration to handle complex tasks, with notable examples like AppAgentX (Jiang et al., 2025) and Mobile-Agent-v2 (Wang et al., 2024a). The latter, upon whose architecture our work builds, employs a framework with specialized planning and reflection agents to overcome long-context challenges. In parallel, training-based methods enhance agent capabilities (Chu et al., 2023; Hong et al., 2024; Lu et al., 2024; Cheng et al., 2025b; Bai et al., 2024; Liu et al., 2025a). For instance, CogAgent (Hong et al., 2024) and UI-TARS (Qin et al., 2025) fine-tune VLMs fine-tunes a VLM for superior GUI grounding, while Reinforcement Learning (RL) has emerged as a dominant paradigm in the GUI agent community. Represented by DigiRL (Bai et al., 2024), this direction has flourished with recent works like MobileRL (Xu et al., 2025), UI-S1 (Lu et al., 2025), and GUI-G2 (Tang et al., 2025) employing diverse optimization strategies. Furthermore, a new wave of reasoning-oriented RL approaches, such as UI-R1 (Lu et al., 2025), GUI-R1 (Luo et al., 2025), and InfiGUI-R1 (Liu et al., 2025c), demonstrates the community’s shift towards agents capable of self-improving through environment feedback. Despite these advances, most efforts emphasize performance and generalization, while the unique safety challenges in mobile phones, driven by dynamic environments with pop-ups, notifications, and inter-app interactions, remain underexplored.
2.2 Adversarial Vulnerabilities in Multimodal Agents
Recent studies reveal that multimodal and LLM-based agents remain broadly vulnerable to adversarial manipulations across both visual and semantic channels. Wu et al. (2024) show that even minor visual perturbations can reliably misguide multimodal agents’ decision pathways. Aichberger et al. (2025) demonstrate that small, human-imperceptible OS-level patches can hijack GUI agents without affecting interface usability. Beyond visual cues, Wang et al. (2025a) introduce cross-modal prompt injections where crafted images override text-based reasoning and safety constraints. In the language domain, Fu et al. (2024) present Imprompter, revealing that structured linguistic sequences can coerce agents into unsafe or unintended tool operations. Cheng et al. (2025a) identify a training-stage backdoor mechanism in which agents are conditioned to trigger harmful actions under particular historical states, highlighting that adversarial risks extend beyond inference into the full development pipeline. These works collectively underscore the inherent fragility of the underlying models driving GUI agents.
@@ 2.3 Environmental Injection Attacks
Within the specific context of GUI agents, environmental dynamics serve as one viable pathway for adversarial risks. The environmental dynamics is prevalent in GUI agents on both smartphones and computers, where notifications, pop-ups, and advertisements continually interfere decision contexts. Ma et al. (2025) highlight that multimodal agents are inherently susceptible to such environmental distractions, where visual noise can easily derail the reasoning process. In this vein, Zhang et al. (2024) demonstrate that adversarial pop-ups can significantly disrupt the task execution of VLM agents in environments such as OSWorld (Xie et al., 2024) and VisualWebArena (Koh et al., 2024), effectively circumventing common defensive prompts and revealing critical vulnerabilities in agent architectures. RiOSWorld (Yang et al., 2025a) categorizes such phenomena as environmental risks and introduces evaluators for intention and completion under dynamic threats. AgentHazard (Liu et al., 2025b) further shows that scalable GUI-hijacking via third-party content can systematically mislead mobile agents. More recently, Chen et al. define the concept of Active Environmental Injection Attacks (AEIA) and demonstrate that adversarial notifications can effectively exploit reasoning gaps in AndroidWorld (Chen et al., 2025). These attacks are often perceptually conspicuous yet strategically deceptive, specifically designed to divert agent behavior through seemingly imperative cues. While these studies often evaluate an agent’s robustness against specific, single injection risks in isolation, they have firmly established the feasibility and severity of dynamic injection attacks, highlighting the significant challenge they pose for comprehensively evaluating the safety of GUI agents.
@@ 2.4 Benchmarks for GUI Agent Security Risk Assessment
The evaluation of GUI agent security is rapidly maturing, with benchmarks evolving to assess risks beyond simple task failure. These can be broadly divided into those targeting specific security vulnerabilities and those measuring overall reliability and policy adherence. For assessing vulnerabilities, benchmarks probe distinct attack vectors. InjecAgent (Zhan et al., 2024) measures an agent’s weakness to indirect prompt injections via tools, while web-focused benchmarks like AdvWeb (Xu et al., 2024) use hidden adversarial prompts to hijack agent actions. In parallel, a second line of work evaluates holistic safety. AgentHarm (Andriushchenko et al., 2024) quantifies the likelihood of an agent to perform harmful actions, and ST-WebAgentBench (Levy et al., 2024) evaluates whether web agents can operate within predefined safety policies. The unique risks on mobile platforms are addressed by dedicated evaluations; MobileSafetyBench (Lee et al., 2024) assesses risk management and VeriOS-Bench (Wu et al., 2025) focuses on verifying agent compliance with system policies, while attack demonstrations like Chen et al. (2025) highlight vulnerabilities to environmental injections via malicious notifications, underscoring the need for such focused security benchmarks.
@@ 3.2 Benchmark Construction and Composition
To realize our dynamic attacks, we implemented a robust, hook-based trigger mechanism that supports evaluating diverse agent frameworks. A predefined agent action (e.g., launching a specific application) activates a hook that broadcasts an adb command. This command is intercepted by a custom on-device helper application, which then renders the adversarial UI element in real-time. This allows us to precisely time our injections. For instance, a deceptive Overlay is presented the moment an agent is about to enter sensitive data. For scenarios involving web-based threats, the helper app can redirect the agent’s browser to a locally hosted, meticulously crafted phishing website that mimics a legitimate service, testing the agent’s ability to discern fraudulent web pages from authentic ones.
@@ 1. RQ1
    How do mobile agents across proprietary, open-source, and fine-tuned models perform on our benchmark in terms of task completion and robustness to attacks?
@@ 4.1 Experimental Setup
We evaluate a representative set of modern Vision-Language Models (VLMs), together with variants across agent frameworks and domain-specific fine-tuning. This set includes leading proprietary models and a representative large-scale open-source model. This includes leading proprietary models and a representative large-scale open-source model. The proprietary systems are: OpenAI’s GPT series (e.g., GPT-4o) (Hurst et al., 2024), Anthropic’s Claude series (e.g., Claude 3.7 Sonnet (Anthropic, 2025)), Google’s Gemini-2.5 Pro (Comanici et al., 2025), and Alibaba’s Qwen-VL-Max (Bai et al., 2023). As a powerful open-source counterpart, we include Qwen2.5-VL-72B-Instruct from the Qwen series (Bai et al., 2025). Detailed model specifications are in Table 7. All VLMs are deployed as mobile GUI agents via two agentic framework, Mobile-Agent v2 (Wang et al., 2024a) and AppAgent (Zhang et al., 2025). For GPT-4o and Qwen2.5-VL-72B-Instruct, we instantiate both controllers to obtain paired comparisons and make the effect of the agent framework on robustness explicit, while the remaining models are evaluated under Mobile-Agent v2. In addition, we include two specialized GUI-oriented VLMs, UI-TARS series (Qin et al., 2025), to assess how domain-specific fine-tuning influences agent reliability under environmental injection attacks. Detailed model and framework specifications are summarized in Table 7.
We evaluate the models’ performance on GhostEI-Bench using a Judge Model approach. The primary metrics include Full Attack Success (FAS), Partial Attack Success (PAS), and Task Completion (TC), as described in Section 3.3.
@@ 4.2 Overall Performance and Vulnerability Analysis
Table 1: Performance of VLMs across different agent frameworks (Mobile-Agent-v2 / AppAgent) and specialized models (UI-TARS). We report absolute counts and percentages for Task Completion (TC), Full Attack Success (FAS), Partial Attack Success (PAS), and Benign Failure (BF). The Vulnerability Rate (VR) is calculated on all non-benign failure cases as defined in Sec $3.$ Higher TC is better; lower FAS, PAS, BF, and VR are better. Bold indicates the best performance in each column.
@@
AppAgent
GPT-4o 37 (33.6%) 24 (21.8%) 12 (10.9%) 38 (34.5%) 50.00
Qwen2.5-VL-72B-Instruct 38 (34.6%) 27 (24.5%) 14 (12.7%) 32 (29.1%) 52.56
UI-TARS
UI-TARS-7B-SFT 29 (26.4%) 20 (18.2%) 16 (14.5%) 46 (41.8%) 56.25
UI-TARS-1.5-7B 45 (40.9%) 19 (17.3%) 20 (18.2%) 27 (24.5%) 46.99
@@
Impact of Agent Frameworks. Our expanded evaluation reveals that the impact of agent frameworks is complex and model-dependent rather than universally beneficial. As shown in Table 1, employing the Mobile-Agent-v2 framework substantially improves the robustness of Qwen2.5-VL (reducing VR from 52.56% to 41.42%) compared to AppAgent. However, for GPT-4o, the same Mobile-Agent-v2 framework actually increases vulnerability (VR: 54.87%) compared to the simpler AppAgent (VR: 50.00%). This contrast demonstrates that different planning frameworks do not inherently guarantee security; instead they introduce specific biases that may either mitigate or amplify vulnerabilities.
Performance of Specialized GUI Fine-Tuning. We further extend our analysis to domain-specific agents by evaluating the UI-TARS series. The results highlight the effectiveness of Supervised Fine-Tuning (SFT) in enhancing GUI grounding. The advanced UI-TARS-1.5-7B achieves a competitive Task Completion rate of 40.9%, significantly outperforming its predecessor and rivaling much larger proprietary models. Interestingly, fine-tuning induces a distinct behavioral shift: these agents tend to remain highly task-oriented even under attack. This point results in lower Full Attack Success rates (17.3%) compared to general VLMs, as the agents are less likely to be fully derailed into a malicious workflow. However, this persistence often leads to higher Partial Attack Success (18.2%), where the agent interacts with deceptive elements while attempting to maintain the original trajectory. This suggests that while SFT successfully enhances execution stability, incorporating explicit safety alignment against dynamic injections represents a promising direction for future optimization.
@@ 4.4 Effects of Reflection and Reasoning Mechanisms
GPT-5-chat-latest (no refl.) 50 (45.5%) 30 (27.3%) 5 (4.6%) 26 (23.6%) 41.67
GPT-5-chat-latest (+refl.) 52 (47.3%) 21 (19.1%) 6 (5.5%) 32 (29.1%) 34.62
@@
Claude-3.7-Sonnet (base) 37 (33.6%) 30 (27.3%) 13 (11.8%) 32 (29.1%) 55.12
Claude-3.7-Sonnet (thinking) 32 (29.1%) 30 (27.3%) 21 (19.1%) 25 (22.7%) 60.0
Reflection Helps Selectively. As shown in Table 2, reflection mechanisms generally reduce vulnerability.This indicates that reflection enables stronger detection of manipulative contexts. GPT-5-chat-latest with reflection also shows moderate gains. In contrast, GPT-4o’s reflection variant improves VR but at the cost of raised benign failure, revealing a trade-off where reflection can sometimes make agents overly cautious.
@@ Appendix C Experiment Settings
Qwen2.5-VL-72B-Instruct API / weights - Qwen Team, Alibaba
UI-TARS-7B-SFT Weights - ByteDance Seed
UI-TARS-1.5-7B Weights -
Appendix D Supplementary Experimental Analysis
In this section, we provide additional experimental results to validate the reliability of our evaluation methodology and explore potential defense mechanisms.
D.1 Reliability Validation of the LLM Judge
The reliability of the automated LLM Judge is paramount to the validity of our benchmark results. Consistent with prevalent methodologies in the GUI agent community (e.g., DigiRL (Bai et al., 2024), DistRL (Wang et al., 2024c)), we employ the LLM-as-a-judge paradigm. To quantitatively validate this approach, we conducted a Human–LLM Agreement study on execution trajectories generated by GPT-4o in the Mobile-Agent-v2 framework.
A human expert independently re-annotated all execution trajectories generated by GPT-4o under the Mobile-Agent v2 framework, and we compared these labels against the original automated GPT-4o evaluations. Among the 110 cases examined, only 7 showed inconsistencies, resulting in a consistency rate of 93.6%. Table 8 presents the full distribution across all five metrics (TC/FAS/PAS/BF/VR).
Table 8: Comparison between GPT-4o Judge and Human Expert annotations.
Judge TC FAS PAS BF VR (%)
GPT-4o (Automated) 38 (34.5%) 33 (30.0%) 12 (10.9%) 28 (25.5%) 54.88
Human Expert 35 (31.8%) 34 (30.9%) 10 (9.9%) 32 (29.1%) 56.41
Difference -2.7% +0.9% -1.0% +3.6% +1.53
We find a strong alignment between the automated GPT-4o judge and the human expert, with a consistency rate of 93.6% in all cases. The few discrepancies stem primarily from ambiguous boundaries in determining whether user instructions should be classified as harmful, rather than from model hallucinations. This high agreement provides empirical evidence that our LLM-as-a-judge protocol is sufficiently reliable for large-scale security evaluation, aligning with methodologies widely adopted in recent GUI-agent research.
D.2 Exploration of Defense Mechanisms: Cautious Prompting
To explore defense mechanisms suggested during peer review, we further evaluate Cautious Prompting (CP), a simple textual intervention that explicitly instructs the agent to prioritize stability and safety awareness. We measure its effect on GPT-4o using the same Mobile-Agent v2 framework. Results are shown in Table 9.
Table 9: Impact of Cautious Prompting (CP) on GPT-4o performance.
Model TC $\uparrow$ FAS $\downarrow$ PAS $\downarrow$ BF $\downarrow$ VR (%) $\downarrow$
GPT-4o (w/o CP) 38 (34.5%) 33 (30.0%) 12 (10.9%) 28 (25.5%) 54.88
GPT-4o (w/ CP) 42 (38.2%) 21 (19.1%) 17 (15.5%) 30 (27.3%) 47.50
Cautious Prompting yields a measurable improvement in robustness, reducing the Vulnerability Rate from 54.88% to 47.50%. This benefit is mainly driven by a substantial reduction in Full Attack Success (from 30.0% to 19.1%), indicating that agents become less likely to execute fully malicious workflows. The increase in Partial Attack Success (10.9% to 15.5%) suggests a shift toward more hesitant or interrupted behaviors rather than complete compliance. Interestingly, Task Completion improves slightly, showing that adding safety constraints does not compromise utility. Overall, these results highlight both the potential and the limitations of prompt-level defenses: while helpful, textual interventions alone cannot fully mitigate visually induced deception.
D.3 Statistical Significance Analysis
To ensure the robustness of our findings against potential execution noise (e.g., network latency, emulator rendering variations), we conducted repeated evaluations. We selected two representative models, Gemini-2.5 Pro and GPT-4o, and performed 3 independent runs for each to quantify the variance in our core metrics.
Table 10: Statistical variance analysis over 3 independent runs (Mean $\pm$ Standard Deviation).
Model TC (%) FAS (%) PAS (%) BF (%) VR (%)
Gemini-2.5 Pro 47.90 $\pm$ 1.85 25.17 $\pm$ 0.47 7.30 $\pm$ 0.73 20.30 $\pm$ 1.85 40.70 $\pm$ 0.50
GPT-4o 34.53 $\pm$ 0.78 30.30 $\pm$ 1.85 10.00 $\pm$ 0.73 25.76 $\pm$ 1.16 54.25 $\pm$ 2.24
As shown in Table 10, the results demonstrate consistently low standard deviations across all evaluation metrics. This stability confirms that the observed performance gaps between models are statistically significant and robust to random execution noise, thereby validating the reliability and reproducibility of the GhostEI-Bench evaluation protocol.
Appendix E Detailed Prompt